Do not calculate the solution to this problem but show how it would be set up. Assume you are trying to create a gel for gel electrophoresis that had a 1.2% concentration of agarose. If you had 50X TBE as the liquid buffer and the gel size was 20mL, how much 50X TBE would you need? How much water?

Receptors in the medulla of the brain will stimulate respiration when hydrogen ion concentration increases. Normal blood pH is 7.4. If Juanita’s blood pH is 6.8, will she be breathing faster or slower than normal? Why?

a.If atoms share an electron pair to become stable, what type of chemical bond is this?

b. Compare covalent and ionic bonds in water.
Answer #1

1) we have to make 20 mL to 1X TBE buffer from 50X TBE buffer

so M1=50X M2=1X V2=20mL V1=?

M1V1=M2V2

V1=M2V2/M1

= 1*20/50

=0.4 mL

1 mL=1000 uL '

so volume of TBE need is 400 uL

, so take 400 uL of TBE and mix it with 19.6 mL (20-0.4=19.6 mL) of water

measuring agarose, pwecentage of agarose in the gel is 1.2%

Mass /volume*100=1.2%

mass/20*100=1.2

mass=1.2*20/100

= 0.24 g

so add 0.24 g of agarose to the prepared TBE solution then melt it , the add EtBr once the mixture cooldowns then pour it into the castor

2) pH=-log [H+]

as the H+ ion concentration increase pH of the solution increases here the pH of the blood has decreased from 7.4 to 6.8 so the receptors in the medulla of the brain will stimulate respiration and rate of respiration increases

3) a) covalent bond is formed by the sharing of electrons it is one the strongest bonds,

b) water do not decrease the strength of covalent bonds but the strength of ionic bond decreases in water because the dielectric constant of water is high, strength, if the ionic bond is inversely proportional to the dielectric constant, dielectric constant of water, is 80 and dielectric constant of air is 1 so in water the strength of ionic bond decreases 80 times. so the strength of ionic bond decrease in water but the strength of covalent bonds is not affected in water.